# Extended-Krylov-subspace methods for trust-region and norm-regularization subproblems

### v2

@@ Abstract
We consider an effective new method for solving trust-region and norm-regularization problems that arise as subproblems in many optimization applications. We show that the solutions to such subproblems lie on a manifold of approximately very low rank as a function of their controlling parameters (trust-region radius or regularization weight). Based on this, we build a basis for this manifold using an efficient extended-Krylov-subspace iteration that involves a single matrix factorization. The problems within the subspace using such a basis may be solved at very low cost using effective high-order root-finding methods. This then provides an alternative to common methods using multiple factorizations or standard Krylov subspaces. We provide numerical results to illustrate the effectiveness of our TREK/NREK approach.
@@ 1 Introduction
We refer to the first possibility as the trivial—or interior—case, while the second is the general one; there is an additional probability-zero possibility, commonly called the hard case—this occurs if $b$ is orthogonal to all the eigenvectors for the leftmost eigenvalue $\lambda_{1}(A)$ and the radius is too large—that we shall ignore for the time being. If we can rule out the trivial and hard cases, (1.5) then suggests an iteration in which, for a sequence of positive $\sigma$, $x(\sigma)$ is found by solving the linear system (1.2), and its suitability assessed by examining the size of the residual $\|x(\sigma)\|-\Delta$.
An interesting question is, therefore, what we can say about the manifold ${\cal S}(A,b)=\mbox{span}\{x(\sigma)\}$ as $\sigma$ varies between $\sigma_{\min}$ and $\infty$. There is no loss of generality, from a theoretical perspective, in considering a problem with a diagonal Hessian, since if $A$ has the spectral factorization
@@
(1.7) ${\begin{array}[t]{c}\mbox{arg min}\vskip-2.84526pt\\ \mbox{ $\scriptstyle y\in\mbox{\raisebox{-0.6pt}{\scriptsize I\thinspace R\thinspace}}^{k}n$ }\end{array}\;}{\scriptstyle\frac{1}{2}}y^{T}\Lambda y-d^{T}y\;\;\mbox{subject to}\;\;\|y\|\leq\Delta$}
@@
for (1.7) are $y_{i}(\sigma)=1/(\lambda_{i}+\sigma)$. If we approximate the solution manifold ${\cal S}(\Lambda,d)$ by the discrete set $\{y(\sigma_{j}),j=1,\ldots,p\}$ at the evaluation points, and compute the singular value decomposition of the resulting $n$ by $p$ matrix $S$ with entries
(1.8) $s_{i,j}=d_{i}/(\lambda_{i}+\sigma_{j})\;\;\mbox{for}\;\;i=1,\ldots,n\;\;\mbox{and}\;\;j=1,\ldots,p,$}
@@
Table 1.1: Singular values of dominant manifold subspace for four eigenvalue distributions
It is then remarkable that the manifold appears to be spanned by so few significant vectors, that is, it is approximately of very low rank. This turns out not to be a coincidence, and we shall exploit this throughout.
This paper is organised as follows. In Section 2, we explain why the solution manifold of interest is (approximately) low rank. We follow this in Section 3 by considering methods that build approximations to this manifold. A particularly appealing approach may be constructed from a so-called extended Krylov subspace, and we show how in Section 4. This then provides the basis for methods aimed at the trust-region subproblem (1.1), and we provide details for positive-definite $A$ in Section 5. We follow this with a description of how we may extend this for general $A$ in Section 6. Numerical experiments follow in Section 7. We show that the same ideas trivially extend to the regularized-norm problem in Section 8, and finally we conclude in Section 9.
@@ 2 Cauchy matrices and accurate low-rank approximation
To investigate the rank of $S$, clearly rows $i$ and $j$ of $S$ are linearly dependent if $\lambda_{i}=\lambda_{j}$, and thus $\mbox{rank}(S)=\mbox{rank}(\bar{S})$, where $\bar{S}$ comprises the rows of $S$ in which each set of such dependencies is replaced by a single representative.
@@
3 Manifold approximation from Krylov subspaces
@@
(3.1) $y_{\ell}={\begin{array}[t]{c}\mbox{arg min}\vskip-2.84526pt\\ \mbox{ $\scriptstyle y\in\mbox{\raisebox{-0.6pt}{\scriptsize I\thinspace R\thinspace}}^{k}$ }\end{array}\;}{\scriptstyle\frac{1}{2}}y^{T}P_{\ell}y-b_{\ell}^{T}y\;\;\mbox{subject to}\;\;\|y\|\leq\Delta.$}
@@
where $x_{\ell}=V_{\ell}y_{\ell}$ and $\|x_{\ell}\|=\Delta$, since $V_{\ell}$ has orthogonal columns (i.e., $V_{\ell}^{T}V_{\ell}=I$). As the solution we seek satisfies $(A+\sigma I)x=b$, and as $V_{\ell}$ will be an increasingly better approximation to the low-dimensional subspace that contains the required solution, it suffices to monitor the residual $r_{\ell}=(A+\sigma_{\ell}I)x_{\ell}-b$ and to stop when $\|r_{\ell}\|$ is sufficiently small. This, however, does not avoid the cost of multiple factorizations; rival methods, e.g., [GoulRobiThor10:mpc, MoreSore83:sisc], also use factorizations, but experience suggests that relatively few often suffice. Another approach for constructing a good approximation of the manifold is to use a method based on a rational Krylov subspace [Ruhe84:LAA, AldaPali25:ARXIV]. However, this will still require solving $\ell$ shifted linear systems.
The alternative is to try to build a good approximation to the manifold by other means. We shall resort to methods based on the extended Krylov subspace (see Section 4).
@@
$y_{i}(\sigma)=\frac{d_{i}}{\lambda_{i}+\sigma},\;\;i=1,\ldots,n,\;\;\mbox{and}\;\;d=U^{T}b.$}
@@ 4 The EKS algorithm
Fortunately, much of the groundwork is already in place. We suppose, for the time being, that $A$ is positive definite. Then Jagels and Reichel [^2] They also propose a second, more expensive algorithm that copes with indefinite $A$. [JageReic09:laa] have shown how to build such a basis if the order [^3] We shall say why our order is more convenient later. in which the components $A^{-m}b$ and $A^{m}b$ arise is interchanged, that is if instead
@@
Using the same logic as in [JageReic09:laa, Alg.2.1], that applies to ${\cal K}^{\prime}_{2m}(A,b)$, it is straightforward to derive the following algorithm, Algorithm 1 1, that instead uses the order ${\cal K}_{2m}(A,b)$.
@@ 16 $\beta_{k}:=u^{T}v_{k}$, $u\leftarrow u-\beta_{k}v_{k}$;
that play the roles of $A_{k}$ in (3.1). It follows from (4.3) that
@@
using (4.5). Again, the orthogonality of the $v_{i}$ and symmetry implies that [^4] When $\delta_{-k-1}=0$, the last three terms in (4.9) are $p_{2k+1,2k+1}=(1-\beta_{-k}\delta_{k})/\beta_{k}$ and $p_{2k+2,2k+1}=p_{2k+3,2k+1}=0$, and thus do not require that we calculate $\alpha_{k}$, $\alpha_{-k-1}$ and $\delta_{k+1}$.
@@
by comparing $p_{2k-1,2k+1}$ with $p_{2k-3,2k-1}$ for subsequent $k$. Finally, if follows from (4.2), (4.5) with $k=1$ and the orthogonality of the $v_{i}$, that [^5] When $\delta_{-1}=0$, $p_{1,1}=1/\beta_{0}$ and $p_{1,2}=p_{1,3}=0$, and thus $\alpha_{0}$, $\alpha_{-1}$ and $\delta_{1}$ need not be calculated.
@@ 5 Solving the convex trust-region subproblem
Note that in the exceptional case where $\delta_{i}=0$, the corresponding $\|r_{i}\|=0$, and thus termination will always occur at this step.
@@ 5.1 Resolves
The need to resolve a problem with the same $A$ and $b$, but a smaller $\Delta$ is a common occurrence in trust-region methods, since such a mechanism is used to guarantee convergence when the current solution has proved unsatisfactory [ConnGoulToin00]. This is easy to accommodate within Algorithm 2; the data generated by the previous call to the algorithm is retained, and the ‘”for loop” re-entered with $k$ taking the value $k_{\mbox{\scriptsize stop}}$ that it last had. The motivation is simply that although the segment of the manifold ${\cal S}(A,b)$ we have observed may now have to expand slightly, it may well be that there is little actual difference; the data provided in Appendix C indicates that in only two of ninety six cases examined were any further extended Krylov iterations necessary, and thus that resolves are generally very inexpensive. Examining the bound (A.6) given in Proposition A.3 in Appendix A, and looking at Figure A.1, we see that the convergence rate of the extended Krylov subspace improves for $\sigma$ beyond a certain point, which is consistent with this observation. Although other methods can also take advantage of resolves, all that we are aware of still require significant extra computation.
@@ 5.2 Other elliptical norms
(5.6) ${\begin{array}[t]{c}\mbox{minimize}\vskip-2.84526pt\\ \mbox{ $\scriptstyle x\in\mbox{\raisebox{-0.8pt}{\footnotesize I\thinspace R\thinspace}}^{n}$ }\end{array}\;}{\scriptstyle\frac{1}{2}}x^{T}Ax-b^{T}x\;\;\mbox{subject to}\;\;\|Px\|\leq\Delta,$}
where $P$ is an easily-invertible matrix—a diagonal $P$ is often a simple but effective choice. On writing
(5.7) $x_{\mbox{\tiny P}}=Px,$}
the minimizer of (5.6) is formally $x_{*}=P^{-1}x_{\mbox{\tiny P}}{}_{*}$, where
(5.8) $x_{\mbox{\tiny P}}{}_{*}={\begin{array}[t]{c}\mbox{arg min}\vskip-2.84526pt\\ \mbox{ $\scriptstyle x\in\mbox{\raisebox{-0.8pt}{\footnotesize I\thinspace R\thinspace}}^{n}$ }\end{array}\;}{\scriptstyle\frac{1}{2}}x_{\mbox{\tiny P}}{}^{T}A_{\mbox{\tiny P}}x_{\mbox{\tiny P}}-b_{\mbox{\tiny P}}{}^{T}x_{\mbox{\tiny P}}\;\;\mbox{subject to}\;\;\|x_{\mbox{\tiny P}}\|\leq\Delta$}
and $A_{\mbox{\tiny P}}=P^{-T}AP^{-1}$ and $b_{\mbox{\tiny P}}=P^{-T}b$. Since the TREK algorithm described above merely relies on products with $A_{\mbox{\tiny P}}$ and its inverse $A_{\mbox{\tiny P}}{}^{-1}=PA^{-1}P^{T}$, once again we merely need a factorization of $A$ (to find $A^{-1}v$ for given $v$) so long as means of inverting $P$ and its transpose are available. Commonly a positive-definite scaling matrix $S$ that is (in some sense) close to $A$ is known, and in this case a decomposition, for example a sparse Cholesky factorization, $S=P^{T}P$ gives the appropriate $P$. We provide details of the necessary changes to devise suitable algorithms to solve the resulting problem
@@
where $\|x\|_{S}\stackrel{{\scriptstyle\rm def}}{{=}}\sqrt{x^{T}Sx}$, in Appendix B. We illustrate one such algorithm here as Algorithm 3 3. Note that here the generated vectors $v_{i}$ now form an $S$-orthonormal basis for the extended Krylov subspace.
@@
It is also worth noting that the extended Krylov subspace (4.1) generated by $A_{\mbox{\tiny P}}$ and $b_{\mbox{\tiny P}}$,
(5.10) ${\cal K}_{2m+1}(A_{\mbox{\tiny P}},b_{\mbox{\tiny P}})=\text{span}\{b_{\mbox{\tiny P}},A_{\mbox{\tiny P}}^{-1}b_{\mbox{\tiny P}},A_{\mbox{\tiny P}}b_{\mbox{\tiny P}},\ldots,A_{\mbox{\tiny P}}^{-m}b_{\mbox{\tiny P}},A_{\mbox{\tiny P}}^{m}b_{\mbox{\tiny P}}\},$}
@@
$P^{-T}\text{span}\{b,(SA^{-1})b,(AS^{-1})b,\ldots,(SA^{-m})b,(AS^{-1})^{m}b\},$}
and it is then possible to derive a suitable variant of Algorithm 1 using $S$ and its inverse rather than $P$ and $P^{-1}$.
@@ 6 Solving the general trust-region subproblem
We have in mind that the shift should chosen be so that $A_{\sigma_{\mbox{\tiny S}}}$ is positive definite. The smallest shift for which $A_{\sigma_{\mbox{\tiny S}}}$ is guaranteed to be positive semi-definite is $\sigma_{\min}$ from (1.3) so any value slightly larger than this is possible. But a far cheaper option that does not rely on knowing the spectrum of $A$, once $A$ has been found not to be positive definite, is to use the value
$\sigma_{\mbox{\tiny G}}=-\min_{1\leq i\leq n}\left[a_{ii}-\sum_{1\leq j\neq i\leq n}|a_{i,j}|\right]+\epsilon_{\mbox{\tiny G}}$}
from the Gershgorin theorem, requiring a single pass through the values of $A$; we have found that this with the heuristic value $\epsilon_{\mbox{\tiny G}}=\sqrt{\epsilon_{\mbox{\tiny M}}}\max_{1\leq i,j\leq n}|a_{i,j}|$ where $\epsilon_{\mbox{\tiny M}}$ is the machine precision is effective.
@@
(6.2) $y_{\ell}={\begin{array}[t]{c}\mbox{arg min}\vskip-2.84526pt\\ \mbox{ $\scriptstyle y\in\mbox{\raisebox{-0.6pt}{\scriptsize I\thinspace R\thinspace}}^{k}$ }\end{array}\;}{\scriptstyle\frac{1}{2}}y^{T}(P_{\ell}-\sigma_{\mbox{\tiny S}}I)y-b_{\ell}^{T}y\;\;\mbox{subject to}\;\;\|y\|\leq\Delta.$}

### v3

# arXiv:2511.11135v3[math.NA] 02 Mar 2026

@@ Abstract
We consider an effective new method for solving trust-region and norm-regularization problems that arise as subproblems in many optimization applications. We show that the solutions to such subproblems as a function of their controlling parameters (trust-region radius or regularization weight). Based on this, we build a basis using an efficient extended-Krylov-subspace iteration that involves a single matrix factorization. The problems within the subspace using such a basis may be solved at very low cost using effective high-order root-finding methods. This then provides an alternative to common methods using multiple factorizations or standard Krylov subspaces. We provide numerical results to illustrate the effectiveness of our TREK/NREK approach.
@@ 1 Introduction
We refer to the first possibility as the trivial—or interior—case, while the second is the general one; there is an additional probability-zero possibility, commonly called the hard case—this occurs if $b$ is orthogonal to all the eigenvectors for the leftmost eigenvalue $\lambda_{1}(A)$ and the radius is too large—that we shall ignore for the time being [^1] A few methods are specifically designed to cope with this possibility [BeckVasi18:siopt, CarmDuch20:sirev].. If we can rule out the trivial and hard cases, (1.5) then suggests an iteration in which, for a sequence of positive $\sigma$, $x(\sigma)$ is found by solving the linear system (1.2), and its suitability assessed by examining the size of the residual $\|x(\sigma)\|-\Delta$.
An interesting question is, therefore, what we can say about the manifold ${\cal M}(A,b)=\{x\mid x=x(\sigma)\;\;\mbox{for some}\;\;\sigma\in[\sigma_{\min},\infty)\}$ or more specifically the ${\cal S}(A,b)=\mbox{span}\{x(\sigma)\}$, that contains ${\cal M}(A,b)$, as $\sigma$ varies between $\sigma_{\min}$ and $\infty$. There is no loss of generality, from a theoretical perspective, in considering a problem with a diagonal Hessian, since if $A$ has the spectral factorization
@@
(1.7) ${\scriptstyle\frac{1}{2}}y^{T}\Lambda y-d^{T}y\;\;\mbox{subject to}\;\;\|y\|\leq\Delta$}
@@
for (1.7) are $1/(\lambda_{i}+\sigma)$. If we approximate the ${\cal S}(\Lambda,d)$ by the discrete set $\{y(\sigma_{j}),j=1,\ldots,p\}$ at the evaluation points, and compute the singular value decomposition of the resulting $n$ by $p$ matrix $S$ with entries
(1.8) $(\lambda_{i}+\sigma_{j})\;\;\mbox{for}\;\;i=1,\ldots,n\;\;\mbox{and}\;\;j=1,\ldots,p,$}
@@
Table 1.1: singular values of for four eigenvalue distributions
It is then remarkable that appears to be spanned by so few significant vectors, that is, it is approximately of very low rank. This turns out not to be a coincidence, and we shall exploit this throughout.
This paper is organised as follows. In Section 2, we explain why the of interest is (approximately) low We follow this in Section 3 by considering methods that build approximations to this . A particularly appealing approach may be constructed from a so-called extended Krylov subspace, and we show how in Section 4. This then provides the basis for methods aimed at the trust-region subproblem (1.1), and we provide details for positive-definite $A$ in Section 5. We follow this with a description of how we may extend this for general $A$ in Section 6. Numerical experiments follow in Section 7. We show that the same ideas trivially extend to the regularized-norm problem in Section 8, and finally we conclude in Section 9.
@@ 2 Cauchy matrices and accurate low-rank approximation
3 approximation from Krylov subspaces
@@
(3.1) $y_{\ell}={\scriptstyle\frac{1}{2}}y^{T}P_{\ell}y-b_{\ell}^{T}y\;\;\mbox{subject to}\;\;\|y\|\leq\Delta$}
@@
where $x_{\ell}=V_{\ell}y_{\ell}$ and $\|x_{\ell}\|=\Delta$, since $V_{\ell}$ has orthogonal columns (i.e., $V_{\ell}^{T}V_{\ell}=I$). As the solution we seek satisfies $(A+\sigma I)x=b$, and as $V_{\ell}$ will be an increasingly better approximation to the low-dimensional subspace that contains the required solution, it suffices to monitor the residual $r_{\ell}=(A+\sigma_{\ell}I)x_{\ell}-b$ and to stop when $\|r_{\ell}\|$ is sufficiently small. This, however, does not avoid the cost of multiple factorizations; rival methods, e.g., [GoulRobiThor10:mpc, MoreSore83:sisc], also use factorizations, but experience suggests that relatively few often suffice. Another approach for constructing a good approximation of is to use a method based on a rational Krylov subspace [Ruhe84:LAA, AldaPali25:ARXIV]. However, this will still require solving $\ell$ shifted linear systems.
The alternative is to try to build a good approximation to by other means. We shall resort to methods based on the extended Krylov subspace (see Section 4).
@@
$\frac{}{\lambda_{i}+\sigma},\;\;i=1,\ldots,n,\;\;\mbox{and}\;\;d=U^{T}b.$}
@@ 4 The EKS algorithm
Fortunately, much of the groundwork is already in place. We suppose, for this section and the next, that $A$ is positive definite333. Then Jagels and Reichel [^4] They also propose a second, more expensive algorithm that copes with indefinite $A$, [JageReic09:laa] have shown how to build such a basis if the order [^5] We shall say why our order is more convenient later. in which the components $A^{-m}b$ and $A^{m}b$ arise is interchanged, that is if instead
@@
Using the same logic as in [JageReic09:laa, Alg.2.1], that applies to ${\cal K}^{\prime}_{2m}(A,b)$, it is straightforward to derive the following algorithm, Algorithm 1 1, that instead uses the order ${\cal K}_{2m}(A,b)$; for completeness, we provide details of an alternative justification, that avoids Laurent polynomials, in Appendix B.
@@ 16 $\beta_{k}:=u^{T}v_{k}$, $u\leftarrow u-\beta_{k}v_{k}$;
that play the roles of in (3.1). It follows from (4.3) that
@@
using (4.5). Again, the orthogonality of the $v_{i}$ and symmetry implies that
@@
by comparing $p_{2k-1,2k+1}$ with $p_{2k-3,2k-1}$ for subsequent $k$. Finally, if follows from (4.2), (4.5) with $k=1$ and the orthogonality of the $v_{i}$, that
@@ 5 Solving the convex trust-region subproblem
Note that in the exceptional case where $\delta_{\ell}=0$, the corresponding $\|r_{\ell}\|=0$, and thus termination will always occur at this step;
@@ 5.1 Resolves
The need to resolve a problem with the same $A$ and $b$, but a smaller $\Delta$ is a common occurrence in trust-region methods, since such a mechanism is used to guarantee convergence when the current solution has proved unsatisfactory [ConnGoulToin00]. This is easy to accommodate within Algorithm 2; the data generated by the previous call to the algorithm is retained, and the “for loop” re-entered with $k$ taking the value $k_{\mbox{\scriptsize stop}}$ that it last had. The motivation is simply that although the to expand slightly, it may well be that there is little actual difference; the data provided in Appendix D indicates that in only cases examined were any further extended Krylov iterations necessary, and thus that resolves are generally very inexpensive. Examining the bound (A.6) given in Proposition A.3 in Appendix A, and looking at Figure A.1, we see that the convergence rate of the extended Krylov subspace improves for $\sigma$ beyond a certain point, which is consistent with this observation. Although other methods can also take advantage of resolves, all that we are aware of still require significant extra computation.
@@ 5.2 Other elliptical norms
(5.6) ${\begin{array}[t]{c}\mbox{minimize}\vskip-2.84526pt\\ \mbox{ $\scriptstyle x\in\mbox{\raisebox{-0.8pt}{\footnotesize I\thinspace R\thinspace}}^{n}$ }\end{array}\;}{\scriptstyle\frac{1}{2}}x^{T}Ax-b^{T}x\;\;\mbox{subject to}\;\;\|x\|\leq\Delta,$}
where is an easily-invertible matrix—a diagonal is often a simple but effective choice. On writing
(5.7) $x_{\mbox{\tiny}}=x,$}
the minimizer of (5.6) is formally $x_{*}=^{-1}x_{\mbox{\tiny}}{}_{*}$, where
(5.8) $x_{\mbox{\tiny}}{}_{*}={\begin{array}[t]{c}\mbox{arg min}\vskip-2.84526pt\\ \mbox{ $\scriptstyle x\in\mbox{\raisebox{-0.8pt}{\footnotesize I\thinspace R\thinspace}}^{n}$ }\end{array}\;}{\scriptstyle\frac{1}{2}}x_{\mbox{\tiny}}{}^{T}A_{\mbox{\tiny}}x_{\mbox{\tiny}}-b_{\mbox{\tiny}}{}^{T}x_{\mbox{\tiny}}\;\;\mbox{subject to}\;\;\|x_{\mbox{\tiny}}\|\leq\Delta$}
and $A_{\mbox{\tiny}}=^{-T}A^{-1}$ and $b_{\mbox{\tiny}}=^{-T}b$. Since the TREK algorithm described above merely relies on products with $A_{\mbox{\tiny}}$ and its inverse $A_{\mbox{\tiny}}{}^{-1}=A^{-1}{}^{T}$, once again we merely need a factorization of $A$ (to find $A^{-1}v$ for given $v$) so long as means of inverting and its transpose are available. Commonly a positive-definite scaling matrix $S$ that is (in some sense) close to $A$ is known, and in this case a decomposition, for example a sparse Cholesky factorization, $S=^{T}$ gives the appropriate . We provide details of the necessary changes to devise suitable algorithms to solve the resulting problem
@@
where $\|x\|_{S}\stackrel{{\scriptstyle\rm def}}{{=}}\sqrt{x^{T}Sx}$, in Appendix C. We illustrate one such algorithm here as Algorithm 3 3. Note that here the generated vectors $v_{i}$ now form an $S$-orthonormal basis for the extended Krylov subspace.
@@
It is also worth noting that the extended Krylov subspace (4.1) generated by $A_{\mbox{\tiny}}$ and $b_{\mbox{\tiny}}$,
(5.10) ${\cal K}_{2m+1}(A_{\mbox{\tiny}},b_{\mbox{\tiny}})=\text{span}\{b_{\mbox{\tiny}},A_{\mbox{\tiny}}^{-1}b_{\mbox{\tiny}},A_{\mbox{\tiny}}b_{\mbox{\tiny}},\ldots,A_{\mbox{\tiny}}^{-m}b_{\mbox{\tiny}},A_{\mbox{\tiny}}^{m}b_{\mbox{\tiny}}\},$}
@@
$^{-T}\text{span}\{b,(SA^{-1})b,(AS^{-1})b,\ldots,b,(AS^{-1})^{m}b\},$}
and it is then possible to derive a suitable variant of Algorithm 1 using $S$ and its inverse rather than and -1.
@@ 6 Solving the general trust-region subproblem
We have in mind that the shift should so that $A_{\sigma_{\mbox{\tiny S}}}$ is positive definite. The smallest shift for which $A_{\sigma_{\mbox{\tiny S}}}$ is guaranteed to be positive semi-definite is $\sigma_{\min}$ from (1.3) so any value slightly larger than this is possible. But a far cheaper option that does not rely on knowing the spectrum of $A$, once $A$ has been found not to be positive definite, is to use the value
$\sigma_{\mbox{\tiny G}}=-\min_{1\leq i\leq n}\left[-\sum_{1\leq j\neq i\leq n}||\right]+\epsilon_{\mbox{\tiny G}}$}
from the Gershgorin theorem, requiring a single pass through the values of $A$; we have found that this with the heuristic value $\epsilon_{\mbox{\tiny G}}=\sqrt{\epsilon_{\mbox{\tiny M}}}\max_{1\leq i,j\leq n}|$ where $\epsilon_{\mbox{\tiny M}}$ is the machine precision is effective.
@@
(6.2) $y_{\ell}={\scriptstyle\frac{1}{2}}y^{T}(P_{\ell}-\sigma_{\mbox{\tiny S}}I)y-b_{\ell}^{T}y\;\;\mbox{subject to}\;\;\|y\|\leq\Delta.$}